OpenStudy (anonymous):

Express (x-6)^2 as an equivalent trinomial

OpenStudy (anonymous):

I don't know how to do this -.-

OpenStudy (anonymous):

\((x-6)^{2}\) = \(x^{2} - 2\times6\times x + 6^{2}\) = \(x^{2} - 12x + 36\)

OpenStudy (anonymous):

Because we know \((a-b)^{2}\) = \(a^{2} - 2ab + b^{2}\)